OpenStudy (anonymous):

Change 2x + 5y = -30 into slope intercept form

OpenStudy (anonymous):

Do you know what slope-intercept form looks like?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Basically it is what you get when you solve this equation for y. So you want to get something like: y = mx + b where m and b will be the numbers for the slope and the intercept.

OpenStudy (anonymous):

So start by subtracting 2x from both sides to isolate your y term

OpenStudy (anonymous):

ok i subttracted 2x from both sides and got 5y = -32

OpenStudy (anonymous):

can someone please help e

OpenStudy (anonymous):

I am helping you..

OpenStudy (anonymous):

You should have 5y = -30 - 2x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You cannot subtract 2x from -30 and get -32

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Because you don't know what x is.

OpenStudy (anonymous):

now do i need to subract 30 ?

OpenStudy (anonymous):

to get -2x by itself?

OpenStudy (anonymous):

i meen add 30 *

OpenStudy (anonymous):

y=-2/5-6

OpenStudy (anonymous):

thankyou!